6.5=0(3.6)+1/2(3.6)
1
Expert's answer
2018-01-24T09:42:08-0500
We can prove that:
6.5=0(3.6)^2+1/2 (3.6)^2
0(3.6)^2=0(12.96)=0
1/2 (3.6)^2=1/2 (12.96)=6.48.
We have 2 significant figures in the answer:
0(3.6)^2+1/2 (3.6)^2=0+6.48=6.48
Round to 6.5.